# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ E.6.4 Our inferences about membership are about sequences, not whole books
We emphasize that our claims about membership are made at the level of sequences, not entire books. For models such as DeepSeek v1, Qwen 2.5, and Gemma 2, where training data are undisclosed, we cannot establish whole-book membership. Instead, our extraction results are about specific sequences drawn from Books3, a known pretraining dataset, and it is at this level that we make membership claims. In some cases, our results are suggestive that entire books were included in training, given that we can extract high-probability sequences spanning a full work (e.g., Appendix D.3.39), but we do not generalize from sequences to whole-book membership. We also do not claim that we know the provenance of those extracted sequences. For Llama models we know that the training data included Books3, and thus at least one copy of each sequence from the books we test was included in the training data via Books3. But this is about all we do know with certainty. We do not know very much about these models’ training data details more generally; knowledge of the inclusion of Books3 is a rare exception, not the norm. Books sequences may have been included in the training data via other unknown sources: other book corpora, widely quoted excerpts in web scrapes, etc. The same could be said for these other models, where we do not know if Books3 was included in the training data. Altogether, these observations highlight that membership questions for LLMs are rarely straightforward. This is already widely known in the literature [e.g., 74], with respect to the presence of duplicates in training datasets [18, 60]. This discussion illustrates how these challenges manifest specifically for books, and why membership claims should be scoped carefully. While our measurements are at the level of individual sequences, but that is not necessarily the only level at which membership is of interest. For copyright, membership may be meaningful in different ways at different scopes—an entire book, a chapter, a single sentence, etc. [76].
@@ E.6.5 Absence of extraction does not imply anything about membership
We state this with respect to describing the problem of membership inference in Appendix E.6.1, but it bears repeating: extraction implies membership, but membership does not imply extraction. This is clear in our results; many books in Books3 have no extractable unique text for Llama models, even though those models were trained on Books3. So when we see $p_{\bm{z}}{=}0\%$, we should not conclude that a given sequence (or the source that contains it) was not included in the training data. This is trivially true for Llama models, but more generally true for models with undisclosed training data. For example, Qwen 2.5 may have been trained on Books3; just because it exhibits low extraction probabilities on Books3 books is not evidence that it was not trained on this corpus. Further, $p_{\bm{z}}{=}0\%$ does not necessarily imply anything about how much a particular sequence may have influenced the model or its outputs. Attribution of generations to training data is a different problem, both with respect to technical work in ML and implications for copyright. We do not study these topics in this work. At a very high level, one can see this distinction through our brief results in Appendix F.2. We run our reconstruction procedure on Sandman Slim [69] and Llama 3.1 70B. This model does not memorize this book at all, with respect to our extraction procedure. We run this experiment to show that it fails—that reconstruction is not something that trivially works for any book. Nevertheless, even though we fail to reconstruct any text from the book, it is clear that the generation we produce was influenced by the text (in an informal sense). This type of influence is not something that we capture or attempt to capture—either formally or informally—in our measurements in this paper.
@@ E.6.6 Extractability and copyright
Appendix F Reconstructing Harry Potter with one seed prompt
Given the extent of memorization that we observe for Harry Potter and the Sorcerer’s Stone [106] (as well as other books) and Llama 3.1 70B, it seemed possible that we could reconstruct the entire book near-verbatim using only a single seed prompt of ground-truth text drawn from the book. This result was a natural follow-on to our memorization experiments—based on the results we obtained—but it is not the main focus of our paper. Therefore, this part of our paper reflects only a preliminary investigation into reconstructing large book segments from a model. It is an existence proof that this type of result is possible, not a comprehensive investigation into how to make this type of method work for different texts.
@@ F.1 Reconstruction method
We take a sliding-context approach that we discuss in Section 7. We take an initial seed $s$ and top it up to a minimum of $50$ tokens (if it is less than that). We use beam search to generate $50$ tokens at a time, and use a maximum context window of $3000$ tokens for each generation. This means that we append the $50$ tokens we generate to the existing context, and clip the first $50$ tokens for iterations that exceed the maximum context. (This is most of the iterations; only at the start do we have a context that can be less than $3000$ tokens). We succeeded with $9$ tries in total. (Our first major success occurred in the third run, when we recovered ten chapters—chapters 1-9, and 12—near-verbatim.) The first $3$ runs were debugging and setup (e.g., testing the key idea with greedy decoding and getting a sense of how short a reasonable seed prompt could be). The last $6$ involved changing the size of the number of maximum context tokens (we settled on $3000$) and the number of beam candidates ($8$), as well as minor tweaks to handling EOS tokens.Handling EOS tokens was the only (minor) complication, and it was simple to resolve. EOS tokens tend to be predicted at the ends of chapters; we therefore remove them and manually replace them with "CHAPTER {n+1}", with {n+1} spelled out (e.g., "TWO", "THREE"). Other attempts involved "Chapter" (instead of all caps) and using the number {n+1} (e.g., "2") instead of spelling out the chapter number (e.g., "TWO"). Other solutions could just down-weight the logits for EOS. Occasionally, the model does not predict an EOS at the end of a chapter, which leads to some misalignment later on (e.g., inserting "CHAPTER FOUR" for the fifth chapter). This can cause the model to repeat segments of chapters that have already been produced. The fix here was also simple. We keep track of how many tokens have been generated since the last EOS token, and if that number surpasses 10,00010{,}000, we assume that we have missed the end of a chapter and account for this (i.e., move the chapter counter ahead even though we did not see an EOS). In future runs, we were able to remove this logic entirely. We defer discussion to future work. The exact code that we ran is in Listing 3. Since this code involves beam search, our results should be (in theory) deterministic. Because in practice hardware non-determinism can complicate this, we ran a (non-exhaustive) test by executing the code on two different sets of $4$ GPUs. Both yielded the same exact output. (Though we also spot-checked some of the underlying logits, and there were some slight differences.)
@@
The substantive diff between the version of Harry Potter and the Sorcerer’s Stone [106] that we have from Books3 and our generated output is minimal. We do some very minor text normalization of the ground-truth book. That is, we trim the beginning and end of the ground-truth book to not contain front matter or anything after the end of the last chapter. We remove _ from both documents (used for italics in the Books3 version) and align ellipses to "..." (which appear as " . . ." in the Books3 version.) We then compute various similarity metrics betwen the two documents. The cosine similarity of TF-IDF vectors of each document is an astounding $0.9999$. TF-IDF is a limited metric, as it treats documents as bags of words and thus fails to capture word order. So, as two additional points of comparison, we compute similarity using greedy longest common subsequence matching at both the word and sentence levels (via difflib’s SequenceMatcher). A score of $1$ indicates a perfect match and 0 indicates no match. We obtain a word-level similarity of $0.992$ and a sentence-level similarity (which is more sensitive to formatting differences) of $0.934$. The qualitative differences involve minor formatting changes (that we do not normalize) and localization differences. For example, the Books3 version has British spelling (e.g., "Mummy" instead of "Mommy"). These formatting and localization differences are the large majority of the diff between the two texts. Occasionally, the model also misses single lines, e.g., “There was a loud ’Oooooh!”’ (Figure 9).One of the most interesting outcomes (that we have not investigated in detail) is very occasional apparent hallucinations. For example, in the first chapter, we observe the following diff:
@@
Selecting the top sequence candidate with beam search always led to this particular outcome in Figure 121, regardless of how we configured our method. Given how specific this particular change was, we investigated it further. It turns out that the generated text reflect the ground-truth from editions of the book printed before 2004. In later editions, this line was changed to the one that we do not generate—the one that is in the ground-truth version we have access to in Books3. This change was done for plot continuity reasons, based on an inconsistency that emerged due to plot development concerning Sirius Black in the later books [58]. This line is not actually hallucinated; it is just from an earlier edition of the book than the one we diff with. Of course, this difference still contributes to our quantitative measurements of similarity, even though it is not a meaningful substantive difference.
@@
F.2 Running reconstruction for books with low memorization signal
As a sanity check, we run the same method on Llama 3.1 70B for a book for which we observe effectively no extracted sequences. For Sandman Slim [69] (Appendix D.3.28), we use the first paragraph as the prompt and (including this prompt, shown in blue) produce a $2000$-token document. Note that the generation gets caught in a loop (we mark where the prompt gets generated in blue later on in the generation.) Interestingly, even though we do not generate anything that we can identify as copied from the book, the generation does contain references to “Sandman Slim,” the main character “James Stark” (who is never referred to directly that way in the book, but as “James Butler Hickok Stark”) as well as another character called “Mason Faim.”
@@ F.2 Running reconstruction for books with low memorization signal
We try much longer prompts as well (e.g., the first two pages of the book), and the model does not continue with text from the book. While these tests are not exhaustive, they are meant to illustrate that this is not behavior that we observe for all books with this approach.
@@ Appendix G Discussion of extended results
We only ran experiments for a small fraction of the entire Books3 dataset—running detailed experiments for $50$ books (Appendix D). We took care to sample books from a variety of sources—in copyright, in the public domain, openly licensed. We included popular books, as well as more obscure ones (Table LABEL:app:tab:sliding-books). Overall, very popular books exhibit the most memorization. It seems likely that these books are duplicated on different parts of the Internet; de-duplication is a challenging problem to implement in practice [74], so it is likely that least some duplicate text persists in training datasets for LLMs.
@@
Llama 1 65B and Llama 2 70B exhibit less memorization of Books3 text both in general (with respect to average extraction rate, see Appendix C) and on the specific books we test. (We do not test Llama 3 70B in our average extraction rate experiments.) In general, we observe a pattern that later generations of Llama models memorize more than earlier ones, with respect to average extraction rates (Appendix C) and extraction coverage (Appendix D.4.1). Most books we test exhibited minimal memorization, measured with respect to probabilistic extraction with top-$k$ decoding (T=1T{=}1, k=40k{=}40) and $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). Our results on longer prefixes (Appendix E), however, indicate that there is a lot more memorization to uncover—even if it is harder to extract, where hardness is loosely equated with requiring a longer prefix.Nevertheless, much of the memorization that we do observe for books in our main sliding-window experiments frequently falls into one of a few categories: copyright notices, lists of ordered (page) numbers (a common artifact of converting epub files to txt format, see Appendix E.4.1), publisher addresses, chapter number listings, and author biographies. All of these are types of text that are highly duplicated (partially or exactly) in various sources, both the Internet and across books. And so, extraction of a copyright notice from a given book does not necessarily mean it was memorized from that book; it was likely memorized due to the presence of numerous similar pieces of text in the training data. The same is also true for author biographies, which are printed on websites, not just in books.Another category was text from popular sources that are quoted within books: quotes from the Bible [123, 122, 23, 130], philosophers like John Stuart Mill [139], classics like those by Dante Alighieri [7], text from U.S. government documents [139, 23], famous political speeches [23], public domain books [136], social media posts referenced in numerous news articles [136], and famous quotes posted on blogs and platforms like Goodreads [89]. In many cases, this was the only such text we were able to extract with non-trivial probability from some books using $50$ token prompts. For our negative controls, we painstakingly verify that each piece of text that we extract within a sequence can be found on multiple sources on the Internet. With respect to the books that we tested that are within the scope of the Kadrey et al. v. Meta, Inc. class action suit [70], we were not able to extract much memorized training data. There were notable exceptions (e.g., Coates [23]). It is of course possible that another extraction technique could reveal additional memorization, but we were unable to meaningfully do so for most books using a $50$-token prefixes. It is also important to note that suit was recently decided in favor of the defendants (Meta), and that memorization of training data was only one issue being argued.
@@
We were able to extract relatively enormous amounts of memorized training data in some cases, from both public domain books (e.g.,Fitzgerald [46], Carroll [19], Joyce [68]) and popular in-copyright books of all stripes (e.g., Rowling [106], Martin [85], Sandberg [112], Brown [11], Tolkien [126], Camus [13]). In particular, we intend to investigate this further for Llama 3+ models, which exhibited an order of magnitude more memorization on some books than our worst-case estimates prior to starting this project. It was this enormous extent of memorization that encouraged us to try to reconstruct a book starting from a single seed prompt; we would not have believed such an outcome was possible prior to measuring memorization in Llama 3+ models—certainly not a flagship model from a frontier company, which continues to be downloaded and used by millions of users (Figure 122).

### v4

@@ E.6.4 Our inferences about membership are about sequences, not whole books
We emphasize that our claims about membership are made at the level of sequences, not entire books. For models such as DeepSeek v1, Qwen 2.5, and Gemma 2, where training data are undisclosed, we cannot establish whole-book membership. Instead, our extraction results are about specific sequences drawn from Books3, a known pretraining dataset, and it is at this level that we make membership claims. In some cases, our results are suggestive that entire books were included in training, given that we can extract high-probability sequences spanning a full work (e.g., Appendix D.3.39), but we do not generalize from sequences to whole-book membership.
We also do not claim that we know the provenance of those extracted sequences. For Llama models we know that the training data included Books3, and thus at least one copy of each sequence from the books we test was included in the training data via Books3. But this is about all we do know with certainty. We do not know very much about these models’ training data details more generally; knowledge of the inclusion of Books3 is a rare exception, not the norm. Books sequences may have been included in the training data via other unknown sources: other book corpora, widely quoted excerpts in web scrapes, etc. The same could be said for these other models, where we do not know if Books3 was included in the training data.
Altogether, these observations highlight that membership questions for LLMs are rarely straightforward. This is already widely known in the literature [e.g., lee2022dedup], with respect to the presence of duplicates in training datasets [carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels]. This discussion illustrates how these challenges manifest specifically for books, and why membership claims should be scoped carefully. While our measurements are at the level of individual sequences, but that is not necessarily the only level at which membership is of interest. For copyright, membership may be meaningful in different ways at different scopes—an entire book, a chapter, a single sentence, etc. [lee2023talkin].
@@ E.6.5 Absence of extraction does not imply anything about membership
We state this with respect to describing the problem of membership inference in Appendix E.6.1, but it bears repeating: extraction implies membership, but membership does not imply extraction. This is clear in our results; many books in Books3 have no extractable unique text for Llama models, even though those models were trained on Books3. So when we see $p_{\bm{z}}{=}0\%$, we should not conclude that a given sequence (or the source that contains it) was not included in the training data. This is trivially true for Llama models, but more generally true for models with undisclosed training data. For example, Qwen 2.5 may have been trained on Books3; just because it exhibits low extraction probabilities on Books3 books is not evidence that it was not trained on this corpus.
Further, $p_{\bm{z}}{=}0\%$ does not necessarily imply anything about how much a particular sequence may have influenced the model or its outputs. Attribution of generations to training data is a different problem, both with respect to technical work in ML and implications for copyright. We do not study these topics in this work. At a very high level, one can see this distinction through our brief results in Appendix F.3. We run our reconstruction procedure on Sandman Slim [sandmanslim] and Llama 3.1 70B. This model does not memorize this book at all, with respect to our extraction procedure. We run this experiment to show that it fails—that reconstruction is not something that trivially works for any book. Nevertheless, even though we fail to reconstruct any text from the book, it is clear that the generation we produce was influenced by the text (in an informal sense). This type of influence is not something that we capture or attempt to capture—either formally or informally—in our measurements in this paper.
@@ E.6.6 Extractability and copyright
Appendix F Reconstructing books with one seed prompt
Given the extent of memorization that we observe for Harry Potter and the Sorcerer’s Stone [hp1] (as well as other books) and Llama 3.1 70B, it seemed possible that we could reconstruct the entire book near-verbatim using only a single seed prompt of ground-truth text drawn from the book.
This result was a natural follow-on to our memorization experiments—based on the results we obtained—but it is not the main focus of our paper. Therefore, this part of our paper reflects only a preliminary investigation into reconstructing large book segments from a model. It is an existence proof that this type of result is possible, not a comprehensive investigation into how to make this type of method work for different texts.
@@ F.1 Reconstruction method
We take a sliding-context approach that we discuss in Section 7. We take an initial seed $s$ and top it up to a minimum of $50$ tokens (if it is less than that). We use beam search to generate $50$ tokens at a time, and use a maximum context window of $3000$ tokens for each generation. This means that we append the $50$ tokens we generate to the existing context, and clip the first $50$ tokens for iterations that exceed the maximum context. (This is most of the iterations; only at the start do we have a context that can be less than $3000$ tokens).
We succeeded with $9$ tries in total. (Our first major success occurred in the third run, when we recovered ten chapters—chapters 1-9, and 12—near-verbatim.) The first $3$ runs were debugging and setup (e.g., testing the key idea with greedy decoding and getting a sense of how short a reasonable seed prompt could be). The last $6$ involved changing the size of the number of maximum context tokens (we settled on $3000$) and the number of beam candidates ($8$), as well as minor tweaks to handling EOS tokens.
Handling EOS tokens was the only (minor) complication, and it was simple to resolve. EOS tokens tend to be predicted at the ends of chapters; we therefore remove them and manually replace them with "CHAPTER {n+1}", with {n+1} spelled out (e.g., "TWO", "THREE"). Other attempts involved "Chapter" (instead of all caps) and using the number {n+1} (e.g., "2") instead of spelling out the chapter number (e.g., "TWO"). Other solutions could just down-weight the logits for EOS.
Occasionally, the model does not predict an EOS at the end of a chapter, which leads to some misalignment later on (e.g., inserting "CHAPTER FOUR" for the fifth chapter). This can cause the model to repeat segments of chapters that have already been produced. The fix here was also simple. We keep track of how many tokens have been generated since the last EOS token, and if that number surpasses 10,00010{,}000, we assume that we have missed the end of a chapter and account for this (i.e., move the chapter counter ahead even though we did not see an EOS). In future runs, we were able to remove this logic entirely. We defer discussion to future work.
The exact code that we ran is in Listing 3. Since this code involves beam search, our results should be (in theory) deterministic. Because in practice hardware non-determinism can complicate this, we ran a (non-exhaustive) test by executing the code on two different sets of $4$ GPUs. Both yielded the same exact output. (Though we also spot-checked some of the underlying logits, and there were some slight differences.)
@@
The substantive diff between the version of Harry Potter and the Sorcerer’s Stone [hp1] that we have from Books3 and our generated output is minimal. We do some very minor text normalization of the ground-truth book. That is, we trim the beginning and end of the ground-truth book to not contain front matter or anything after the end of the last chapter. We remove _ from both documents (used for italics in the Books3 version) and align ellipses to "..." (which appear as " . . ." in the Books3 version.)
We then compute various similarity metrics betwen the two documents. The cosine similarity of TF-IDF vectors of each document is an astounding $0.9999$. TF-IDF is a limited metric, as it treats documents as bags of words and thus fails to capture word order. So, as two additional points of comparison, we compute similarity using greedy longest common subsequence (LCS) matching at both the word and sentence levels (via difflib’s SequenceMatcher). A score of $1$ indicates a perfect match and 0 indicates no match. We obtain a word-level similarity of $0.992$ and a sentence-level similarity (which is more sensitive to formatting differences) of $0.934$.
The qualitative differences involve minor formatting changes (that we do not normalize) and localization differences. For example, the Books3 version has British spelling (e.g., "Mummy" instead of "Mommy"). These formatting and localization differences are the large majority of the diff between the two texts. Occasionally, the model also misses single lines, e.g., “There was a loud ’Oooooh!”’ (Figure 9).
One of the most interesting outcomes (that we have not investigated in detail) is very occasional apparent hallucinations. For example, in the first chapter, we observe the following diff:
@@
Selecting the top sequence candidate with beam search always led to this particular outcome in Figure 121, regardless of how we configured our method. Given how specific this particular change was, we investigated it further. It turns out that the generated text reflect the ground-truth from editions of the book printed before 2004. In later editions, this line was changed to the one that we do not generate—the one that is in the ground-truth version we have access to in Books3. This change was done for plot continuity reasons, based on an inconsistency that emerged due to plot development concerning Sirius Black in the later books [sirius]. This line is not actually hallucinated; it is just from an earlier edition of the book than the one we diff with. Of course, this difference still contributes to our quantitative measurements of similarity, even though it is not a meaningful substantive difference.
@@
F.2 Running partial reconstruction for books with partial high memorization signal
We ran the exact same script on the essay “The Case for Reparations,” which is included in Ta Nehisi-Coates’ book, When We Were Eight Years in Power [wewere]. We observed enormous amounts of memorization for this essay in our sliding-window experiments. (See Section 5.1, highly memorized middle section; Appendix E.3). This was one of the books in the Kadrey v. Meta, Inc. [kadreyamendedconsolidated] lawsuit, for which the court in part ruled for fair use because neither party produced evidence that demonstrated sizable amounts of extraction of training data of the books in scope for the suit [kadreyjudgment].
We summarize the first run of this experiment, where we ran the exact same code that reconstructed the first Harry Potter on the first line of the essay ($31$ tokens, including BOS):
  • Clyde Ross was born in 1923, the seventh of thirteen children, near Clarksdale, Mississippi, the home of the blues.
[…]
F.3 Running reconstruction for books with low memorization signal
As a sanity check, we run the same method on Llama 3.1 70B for a book for which we observe effectively no extracted sequences.
For Sandman Slim [sandmanslim] (Appendix D.3.28), we use the first paragraph as the prompt and (including this prompt, shown in blue) produce a $2000$-token document.
Note that the generation gets caught in a loop (we mark where the prompt gets generated in blue later on in the generation.) Interestingly, even though we do not generate anything that we can identify as copied from the book, the generation does contain references to “Sandman Slim,” the main character “James Stark” (who is never referred to directly that way in the book, but as “James Butler Hickok Stark”) as well as another character called “Mason Faim.”
@@ F.3 Running reconstruction for books with low memorization signal
We try much longer prompts as well (e.g., the first two pages of the book), and the model does not continue with text from the book.
While these tests are not exhaustive, they are meant to illustrate that this is not behavior that we observe for all books with this approach.
@@ Appendix G Discussion of extended results
We only ran experiments for a small fraction of the entire Books3 dataset—running detailed experiments for $50$ books (Appendix D). We took care to sample books from a variety of sources—in copyright, in the public domain, openly licensed. We included popular books, as well as more obscure ones (Table LABEL:app:tab:sliding-books). Overall, very popular books exhibit the most memorization. It seems likely that these books are duplicated on different parts of the Internet; de-duplication is a challenging problem to implement in practice [lee2022dedup], so it is likely that least some duplicate text persists in training datasets for LLMs.
@@
Llama 1 65B and Llama 2 70B exhibit less memorization of Books3 text both in general (with respect to average extraction rate, see Appendix C) and on the specific books we test. (We do not test Llama 3 70B in our average extraction rate experiments.) In general, we observe a pattern that later generations of Llama models memorize more than earlier ones, with respect to average extraction rates (Appendix C) and extraction coverage (Appendix D.4.1).
Most books we test exhibited minimal memorization, measured with respect to probabilistic extraction with top-$k$ decoding (T=1T{=}1, k=40k{=}40) and $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). Our results on longer prefixes (Appendix E), however, indicate that there is a lot more memorization to uncover—even if it is harder to extract, where hardness is loosely equated with requiring a longer prefix.
Nevertheless, much of the memorization that we do observe for books in our main sliding-window experiments frequently falls into one of a few categories: copyright notices, lists of ordered (page) numbers (a common artifact of converting epub files to txt format, see Appendix E.4.1), publisher addresses, chapter number listings, and author biographies. All of these are types of text that are highly duplicated (partially or exactly) in various sources, both the Internet and across books. And so, extraction of a copyright notice from a given book does not necessarily mean it was memorized from that book; it was likely memorized due to the presence of numerous similar pieces of text in the training data. The same is also true for author biographies, which are printed on websites, not just in books.
Another category was text from popular sources that are quoted within books: quotes from the Bible [embraced, unglued, wewere, emperorgladness], philosophers like John Stuart Mill [zittrain], classics like those by Dante Alighieri [dante], text from U.S. government documents [zittrain, wewere], famous political speeches [wewere], public domain books [careless], social media posts referenced in numerous news articles [careless], and famous quotes posted on blogs and platforms like Goodreads [beloved]. In many cases, this was the only such text we were able to extract with non-trivial probability from some books using $50$ token prompts. For our negative controls, we painstakingly verify that each piece of text that we extract within a sequence can be found on multiple sources on the Internet.
With respect to the books that we tested that are within the scope of the Kadrey et al. v. Meta, Inc. class action suit [kadreyamendedconsolidated], we were not able to extract much memorized training data. There were notable exceptions (e.g., wewere). It is of course possible that another extraction technique could reveal additional memorization, but we were unable to meaningfully do so for most books using a $50$-token prefixes. It is also important to note that suit was recently decided in favor of the defendants (Meta), and that memorization of training data was only one issue being argued.
@@
We were able to extract relatively enormous amounts of memorized training data in some cases, from both public domain books (e.g.,gatsby, alice, ulysses) and popular in-copyright books of all stripes (e.g., hp1, thrones, leanin, davinci, hobbit, sisyphus). In particular, we intend to investigate this further for Llama 3+ models, which exhibited an order of magnitude more memorization on some books than our worst-case estimates prior to starting this project. It was this enormous extent of memorization that encouraged us to try to reconstruct a book starting from a single seed prompt; we would not have believed such an outcome was possible prior to measuring memorization in Llama 3+ models—certainly not a flagship model from a frontier company, which continues to be downloaded and used by millions of users (Figure 122).